Maple Leaf Production manufactures truck tires. The following information is available for the last operating period.

  • Maple Leaf produced and sold 95,000 tires for $38 each. Budgeted production was 100,000 tires.
  • Standard variable costs per tire follow.
Direct materials: 4 pounds at $2.00 $ 8.00
Direct labor: 0.35 hours at $16.00 5.60
Variable production overhead: 0.15 machine-hours at $15 per hour 2.25
Total variable costs $ 15.85
  • Fixed production overhead costs:

Monthly budget $2,350,000

  • Fixed overhead is applied at the rate of $23.50 per tire.
  • Actual production costs:
Direct materials purchased and used: 399,000 pounds at $1.70 $ 678,300
Direct labor: 30,500 hours at $16.30 497,150
Variable overhead: 15,000 machine-hours at $15.80 per hour 237,000
Fixed overhead 2,360,000

Required:
a.
Prepare a cost variance analysis for each variable cost for Maple Leaf Productions.
b. Prepare a fixed overhead cost variance analysis.
c. (Appendix) Prepare the journal entries to record the activity for the last period using standard costing. Assume that all variances are closed to cost of goods sold at the end of the operating period.
Answer #1
Maple Leaf
a) Direct Material cost Variance=AQ*AP-SQ*SP
AQ*AP=(A) $                               6,78,300.00
SQ*SP=(95000 units produced*4 pound*$2)=(B) $                               7,60,000.00
Direct Material cost Variance=(A)-(B) $                                  81,700.00 (F)
Material Price Variance=AQ*(AP-SP)
AQ*AP=(A) $                               6,78,300.00
AQ*SP=(399000*$2)=(B) $                               7,98,000.00
Material Price Variance=(A)-(B) $                               1,19,700.00 (F)
Material Efficiency Variance=SP*(AQ-SQ)
SP*AQ=($2*399000)=(A) $                               7,98,000.00
SP*SQ=($2*95000*4)=(B) $                               7,60,000.00
Material Efficiency Variance=(A)-(B) $                                  38,000.00 (U)
Direct Labor cost Variance=AH*AR-SH*SR
AH*AR=(A) $                               4,97,150.00
SH*SR=(95000 units*.35*$16)=(B) $                               5,32,000.00
Direct Labor cost Variance=(A)-(B) $                                  34,850.00 (F)
Labor Rate Variance=AH*(AR-SR)
AH*AR=(A) $                               4,97,150.00
AH*SR=(B)=(30500*$16) $                               4,88,000.00
Labor Rate Variance=(A)-(B) $                                     9,150.00 (U)
Labor Efficiency Variance=SR*(AH-SH)
SR*AH=($16*30500)=(A) $                               4,88,000.00
SR*SH=($16*95000*.35)=(B) $                               5,32,000.00
Labor Efficiency Variance=(A)-(B) $                                  44,000.00 (F)
Variable Overhead Variance=AH*AR-SH*SR
AH*AR=(A) $                               2,37,000.00
SH*SR=(95000*.15*$15)=(B) $                               2,13,750.00
Variable Overhead Variance=(A)-(B) $                                  23,250.00 (U)
Variable Overhead Rate Variance=AH*(AR-SR)
AH*AR=(A) $                               2,37,000.00
AH*SR=(B)=(15000*$15) $                               2,25,000.00
Variable Overhead Rate Variance=(A)-(B) $                                  12,000.00 (U)
Variable Overhead Efficiency Variance=SR(AH-SH)
SR*AH=($15*15000)=(A) $                               2,25,000.00
SR*SH=($15*95000*.15) $                               2,13,750.00
Variable Overhead Efficiency Variance=(A)-(B) $                                  11,250.00 (U)
b) Actual Fixed Overhead $                            23,60,000.00
Applied Fixed Overhead=Budgeted Units * Applied Overhead rate=(95000*$23.50) $                            22,32,500.00
Fixed Overhead cost variance $                               1,27,500.00 (U)
c ) Particular Debit Credit
Work in Process Inventory $                               7,60,000.00
Material Efficiency Variance $                                  38,000.00
    To Material Price Variance $              1,19,700.00
    To Accounts Payable $              6,78,300.00
Work in Process Inventory $                               5,32,000.00
Labor Rate Variance $                                     9,150.00
    To Labor Efficiency Variance $                 44,000.00
    To Accounts Payable $              4,97,150.00
Work in Process Inventory $                               2,13,750.00
     To Variable Overhead applied $              2,13,750.00
Variable Overhead actual $                               2,37,000.00
     To Miscellaneous Payable and Inventory accounts $              2,37,000.00
Variable Overhead applied $                               2,13,750.00
Variable Overhead rate Variance $                                  12,000.00
Variable Overhead Efficiency variance $                                  11,250.00
    To Variable Overhead actual $              2,37,000.00
Work in Process Inventory(95000*$23.50) $                            22,32,500.00
   To Fixed Overhead Applied $           22,32,500.00
Actual Fixed Overhead $                            23,60,000.00
    To Miscellaneous Payable and Inventory Accounts $           23,60,000.00
Fixed Overhead applied $                            22,32,500.00
Fixed Overhead Volume Variance(Budgeted -Applied)=($2350000-$2232500) $                               1,17,500.00
Fixed Overhead Price Variance(Actual-Budgeted)=($2360000-2350000) $                                  10,000.00
    To Actual Fixed overhead $           23,60,000.00
Finished Goods Inventory $                            37,38,250.00
    To Work in Process($760000+$532000+$213750+$2232500) $           37,38,250.00
Accounts Receivable(95000*$38) $                            36,10,000.00
    To Sales Revenue $           36,10,000.00
Cost of goods sold $                            37,38,250.00
    To Finished Goods Inventory $           37,38,250.00
Cost of goods sold(Balancing figure) $                                  34,200.00
Material Price Variance $                               1,19,700.00
Labor efficiency variance $                                  44,000.00
    To Material Efficiency Variance $                 38,000.00
    To Labor Rate Variance $                    9,150.00
   To Variable Overhead Rate Variance $                 12,000.00
   To Variable Overhead Efficiency Variance $                 11,250.00
    To Fixed Overhead Volume Variance $              1,17,500.00
    To Fixed Overhead Price Variance $                 10,000.00